# A minimal model for multigroup adaptive SIS epidemics

### v1

@@ Abstract
We propose a generalization of the adaptive N-Intertwined Mean-Field Approximation (aNIMFA) model studied in Achterberg and Sensi [1] to a heterogeneous network of communities. In particular, the multigroup aNIMFA model describes the impact of both local and global disease awareness on the spread of a disease in a network. We obtain results on existence and stability of the equilibria of the system, in terms of the basic reproduction number $R_{0}$. Under light constraints, we show that the basic reproduction number $R_{0}$ is equivalent to the basic reproduction number of the NIMFA model on static networks. Based on numerical simulations, we demonstrate that with just two communities periodic behaviour can occur, which contrasts the case with only a single community, in which periodicity was ruled out analytically. We also find that breaking connections between communities is more fruitful compared to breaking connections within communities to reduce the disease outbreak on dense networks, but both strategies are viable to networks with fewer links. Finally, we emphasise that our method of modelling adaptivity is not limited to SIS models, but has huge potential to be applied in other compartmental models in epidemiology.
@@ 1 Introduction
Various approaches have been suggested to model the response of individuals to an epidemic. A common approach is to apply an SIR-like (Susceptible – Infected – Recovered) compartmental model by including an Aware compartment [2, 3, 4, 5, 6]. Aware individuals are conscious of the ongoing disease and are therefore more cautious towards contact with other individuals. Thus, aware individuals are less likely to be infected. Moreover, some authors included the possibility of corruption of information spreading throughout a population [7, 8, 9], which has been very problematic throughout the COVID-19 pandemic in particular [10].
@@ 2 The multigroup aNIMFA model
where $y_{i}$ is the local prevalence in group $i$, $z_{ij}$ is the link density between group $i$ and $j$, $\delta_{i}$ is the curing rate of group $i$, $\beta_{ij}$ is the infection rate from group $j$ to group $i$, $\zeta_{ij}$ is the link-breaking rate between group $i$ and $j$, $\xi_{ij}$ is the link-creation rate between group $i$ and $j$ and $\bar{y}$ is the global prevalence (or simply the prevalence) of the infectious disease in the whole population, i.e. is the average prevalence over all communities. We are implicitly assuming here that all communities have the same size. In general, the functional responses $f_{\textnormal{br,ij}}$ and $f_{\textnormal{cr,ij}}$ can be different between each region $i$ and $j$. In particular, when $i=j$, the functional responses $f_{\textnormal{br,ii}}$ and $f_{\textnormal{cr,ii}}$ describe the different responses of each community $i$ to the internal spread of the infectious disease.
@@
The $n$ communities can be represented as $n$ nodes in a graph $G$, where each node represents one community. The links in the graph $G$ constitute the connections $z_{ij}$ between community $i$ and $j$. We assume the graph $G$ is connected (i.e. the adjacency matrix $B=(\beta_{ij})$ is irreducible). From a biological point of view, a natural assumption would be to consider all $\delta_{i}$ equal, since they represent the inverse of the recovery rate of the same disease. However, each region may have a different healthcare quality, leading to potentially slightly different recovery rates $\delta_{i}$. Therefore, we keep the model (2) as general as possible.
@@
Equation (2) is coined the multigroup adaptive N-Intertwined Mean-Field Approximation (multigroup aNIMFA), named after the NIMFA model [27]. On a technical note, Eq. (2) contains $n^{2}+n$ equations, not $n$, but we will adhere to the definition (2) due its analogy to the static NIMFA model [27]. We assume the parameters $\delta_{i}$, $\beta_{ij}$, $\zeta_{ij}$, $\xi_{ij}$ can be different for each node and link, and are assumed to be positive. The standard NIMFA model with link weights $z_{ij}=z_{ij}(0)$ is recovered if $\zeta_{ij}=\xi_{ij}=0$ for all $i$ and $j$. Furthermore, $f_{\textnormal{br,ij}}$ and $f_{\textnormal{cr,ij}}$ are assumed to be non-negative for all $0\leq y\leq 1$ and all $i$ and $j$. We emphasise here that the link density in the aNIMFA model is not necessarily symmetric, i.e. $z_{ij}\neq z_{ji}$. This allows us to capture unilateral decisions to limit or encourage movement between community $i$ and community $j$, but not vice versa. For example, we can think of some country A which lifts restrictions on incoming flights from another country B, even though country B still discourages its inhabitants to travel.
@@ 3.3 Basic Reproduction Number
Then, the basic reproduction number $R_{0}$ follows as the largest eigenvalue $\rho$ of the matrix $F$
@@
Assuming no link-breaking at zero prevalence, i.e. $f_{\textnormal{br,ij}}(\mathbf{0})=0$ for all $i,j$, then the steady-state link density $z_{ij}^{\textnormal{DFE}}=1$ for all $i,j$ and the matrix
@@ 3.4 Stability of the DFE
The proof of corollary 4 coincides, up to extremely minor adjustements, with the second half of the proof of [28, Thm. 1], which in turn makes use of a few smaller results presented in the same paper; for the sake of brevity, we do not repeat it here.
@@ 3.5 Endemic Equilibria
We now prove, under slightly more restrictive assumptions than $R_{0}>1$, the existence of (at least one) Endemic Equilibrium (EE).
@@
Our proof of Theorem 5 relies heavily on the assumption on the minimum row/column sum being strictly bigger than 1, hence it fails so for a generic matrix $F$, if we only assume $\rho(F)>1$. However, the proof of Theorem 5 makes use of the Poincaré-Miranda theorem, which might not be necessary to prove this result.
@@ 4 Numerical simulations
The complexity of the governing equations (2) complicates deriving further results on the dynamics. We therefore resort to numerical simulations of system (2) in specific scenarios. The simulations in this section have been executed in Matlab and are based on a simple Forward Euler scheme with time step $\Delta t=0.01$ of system (2). The code is available from the authors upon reasonable request.
We present four cases studies on the multigroup aNIMFA model, each discussing a key aspect of the interplay between infectious disease dynamics and disease awareness.
@@ 4.1 Case study 1: Periodic behaviour
Figure 1: Case study 1: effect of asymmetric response between two communities. Parameters are $n=2$, $\beta_{ij}\sim U([0,1.2]),\delta_{i}=U([0,0.5]),\zeta_{ij}\sim U([0,1]),\xi_{ij}\sim U([0,1])$ for $i,j=1,2$, where $U$ denotes the uniform distribution. (a) local and average prevalence; (b) link densities. We observe a quick convergence towards a stable limit cycle, representing endemicity of the disease alternating between high and low prevalence in the population.
@@
Even though the provided functional responses contain indicator functions, which are not $\mathcal{C}^{1}$-functions over the interval $[0,1]$, we emphasise that other choices of $f_{\text{cr,ii}}$, $f_{\text{br,ii}}$, $f_{\text{cr,ij}}$ and $f_{\text{br,ij}}$ can also lead to periodic behaviour. The benefit of having $\mathcal{C}^{1}$-functions is that the solution is guaranteed to be unique. One method is to approximate the indicator function111 A good approximation of the indicator function of the interval $[a,1]\subset(0,1]$ is $f(x)=\dfrac{1}{1+\left(x^{\frac{\ln(2)}{\ln(a)}}-1\right)^{k}},\quad\hbox{with}\quad k\in 2\mathbb{N},k\gg 1.$ , for which we verified that periodic solutions also occur.
@@
4.2 Case study 2: Seasonality
Besides periodic behaviour emerging naturally from the multigroup aNIMFA model, it is natural for infectious diseases to exhibit seasonal dependency. In particular, the flu is known to resurface in the colder months, because of the reduced functioning of the immune system and the increased crowding inside buildings. The natural periodicity therefore appears through the infection rates $\beta_{ij}$ with period $T$. The collective behaviour of a population usually changes gradually throughout the year and slowly follows the temperature changes. However, modelling abrupt (i.e. discontinuous) changes can be of interest as well, especially considering the possibility of overnight introduction or lifting of lockdown measures, which have been a significant part of the containment strategy of COVID-19 in various nations.
As periodic functions we utilise the sine wave and the block wave with period $T$;
$f(t)$ $=\sin^{2}\left(\frac{\pi t}{T}\right)$ (13a)
$f(t)$ $=4\left\lfloor\frac{t}{T}\right\rfloor-2\left\lfloor\frac{2t}{T}\right\rfloor+1$ (13b)
Consider a complete network of $n=20$ communities, that is, all communities are assumed to be connected to all other communities. A disease emerges emerges in community 1 at time $t=0$, which translates into the initial conditions $y_{1}(0)=0.2$ and $y_{i}(0)=0$ for all other communities $i$. The initial link density $z_{ij}(0)=1$ for all communities $i$. Each community is assumed to have only local awareness, i.e. the link-breaking and link-creation mechanisms are independent of the global prevalence $\bar{y}$. For the link-creation mechanism, we propose
and for the link-breaking mechanism, we pick
$\begin{split}f_{\textnormal{br,ii}}&=y_{i}^{2},\\ f_{\textnormal{br,ij}}&=y_{i}y_{j}.\end{split}$ (15)
We quantify the impact of the periodic functions on the size of the epidemic outbreak based on two parameters: the global prevalence in the steady state $\bar{y}_{\infty}$ and the peak prevalence $y_{p}$, which is the maximal prevalence over all times and all communities:
$y_{p}:=\max_{\begin{subarray}{c}0\leq t\leq\infty\\ i=1,...,n\end{subarray}}y_{i}(t)$
Figure 2 demonstrates the prevalence for each of the periodic infection rates from Eq. (13). Figure 2 depicts that a block wave is more devastating than a sine wave in terms of the peak prevalence. In other words, a sudden uprising and disappearance of a disease is much more dangerous than gradual outbreaks. For sudden uprisings, people cannot respond to the disease sufficiently fast and a large outbreak is inevitable. Therefore, timely communication from governments and health agencies is crucial to mitigate dangerous epidemic outbreaks.
(a) Sine wave
(b) Block wave
Figure 2: Case study 2: effect of seasonality on the prevalence in terms of a time-periodic infection rate with period $T=10$ using (a) a sine wave and (b) a block wave. Parameters are $n=20$, $\beta_{ij}\sim U([0,0.25]),\delta_{i}=1,\zeta_{ij}\sim U([0,2]),\xi_{ij}\sim U([0,1])$ for all $i,j$, where $U$ denotes the uniform distribution. The basic reproduction number $R_{0}=2.53$ for $f(t)=1$.
We further quantify the differences between sine and block waves in terms of the average prevalence over one period and the peak prevalence. Figure 3 depicts the dependence of the properties on the period $T$. The black, dashed baseline in Figure 3 corresponding to $f(t)=0.5$ is equal to the average over one period of the periodic functions, i.e. the total force of infection is assumed to be constant. Figure 3 indeed shows that the block waves are the most disastrous in terms of the peak prevalence, for any period $T$. Moreover, longer time periods result in more severe pandemics in terms of the peak prevalence, because during a long period of high infectiousness, the disease can spread, but the network cannot adapt sufficiently fast, leading to higher prevalence values. On the other hand, the average prevalence over one period appears to decrease due to long periods of inactivity.
Figure 3: Case study 2: impact of the period $T$ on the steady-state prevalence $\bar{y}_{\infty}$ and the peak prevalence $y_{p}$ for sine waves (green) and block waves (red) for the peak prevalence (triangle) and average prevalence (circle). Parameters are the same as Figure 2.
We now investigate the seasonal behaviour for two types of networks: the Barbell graph (two complete graphs connected by a single edge; see Figure 4 for a visualization) and the star graph (see Figure 6 for a visualization).
Figure 4: Barbell graph with 12 nodes.
With the same parameters and choice of periodic functions as in Figure 2, the Barbell graph shows disease persistence with $R_{0}>1$, leading to periodic behavior. For the Barbell graph, Figure 5 illustrates that periodic behavior appears immediately, and changing the response function (i.e., from sine to block wave) results in a similar peak prevalence. However, the wave generated with the block function is peculiar: after the peak, there is an initial slow decline followed by a steeper decline. This distinct behavior can be attributed to the structure of the Barbell graph. The Barbell graph consists of two cliques connected by one or more bridging nodes. Without loss of generality, in our case we consider only one bridge node.
When the disease initially starts in one of the cliques, the high connectivity within the clique causes a rapid initial spread, leading to an overshoot in the initial prevalence. This initial overshoot is because the dense connections within the clique facilitate a swift transmission of the disease, causing the number of infected individuals to rise quickly. As the disease spreads through the bridge node to the second clique, the overall prevalence reaches its peak. The subsequent decline in the number of infected individuals is initially slow due to the high number of connections within each clique, which sustains the transmission for a longer period within the clique before the disease starts to die out. Once the disease prevalence in the first clique significantly reduces, the decline becomes steeper as the disease runs out of susceptible individuals to infect. The second, faster and steeper decline is caused by the switching “off” of the block wave. The abrupt change in the response function results in an exponential convergence towards zero prevalence, i.e., the Disease-Free Equilibrium (DFE). Finally, the subsequent switching “on” before reaching the DFE, which only happens asymptotically as $t\to+\infty$, allows for a rapid increase in infection again.
(a) Sine wave
(b) Block wave
Figure 5: Case study 2: effect of seasonality on the prevalence in terms of a time-periodic infection rate with period $T=10$ using (a) a sine wave and (b) a block wave for a Barbell graph on $n=20$ nodes. The parameters and functional responses are the same as the complete network, which correspond to $R_{0}=9.59$.
Figure 6: Star graph with 7 nodes.
For the star graph (visualized in Figure 6), using the parameters from Section 4.2 results in $R_{0}<1$, consequently leading directly to extinction of the epidemic. By slightly altering the parameters, we find that the star graph in Figure 7 exhibits periodic waves similar to those of the complete graph; however, the peaks in the hub community appear to dominate the peaks of all other nodes. After all, the hub is connected to all other nodes and has therefore the most chance of getting infected.
(a) Sine wave
(b) Block wave
Figure 7: Case study 2: effect of seasonality on the prevalence in terms of a time-periodic infection rate with period $T=10$ using (a) a sine wave and (b) a block wave for the star graph on $n=20$ nodes. The functional responses are the same as in the complete network, with the parameters $\beta_{ij}\sim U([0,1]),\delta_{i}=1,\zeta_{ij}\sim U([0,2]),\xi_{ij}\sim U([0,1])$ for all $i,j$, where $U$ denotes the uniform distribution. The basic reproduction number $R_{0}=2.99$.
4.3 Case study 3: Internal and external connections

### v3

@@ Abstract
We propose a generalization of the adaptive N-Intertwined Mean-Field Approximation (aNIMFA) model studied in Achterberg and Sensi [1] to a heterogeneous network of communities. In particular, the multigroup aNIMFA model describes the impact of both local and global disease awareness on the spread of a disease in a network. We obtain results on existence and stability of the equilibria of the system, in terms of the basic reproduction number $R_{0}$. Assuming individuals have no reason to decrease their contacts in the absence of disease, we show that the basic reproduction number $R_{0}$ is equivalent to the basic reproduction number of the NIMFA model on static networks. Based on numerical simulations, we demonstrate that with just two communities periodic behaviour can occur, which contrasts the case with only a single community, in which periodicity was ruled out analytically. We also find that breaking connections between communities is more fruitful compared to breaking connections within communities to reduce the disease outbreak on dense networks, but both strategies are viable to networks with fewer links. Finally, we emphasize that our method of modelling adaptivity is not limited to SIS models, but has huge potential to be applied in other compartmental models in epidemiology.
@@ 1 Introduction
During epidemic outbreaks without access to medicines, vaccines and other pharmaceutical prevention measures, the only possibility to reduce the spread of the outbreak is by contact avoidance. If the choice for contact avoidance is based on the prevalence of the disease in the population, we call it adaptivity. Various approaches have been suggested to model the response of individuals to an epidemic. A common approach is to apply an SIR-like (Susceptible – Infected – Recovered) compartmental model by including an Aware compartment [2, 3, 4, 5, 6]. Aware individuals are conscious of the ongoing disease and are therefore more cautious towards contact with other individuals. Thus, aware individuals are less likely to be infected. Moreover, some authors included the possibility of corruption of information spreading throughout a population [7, 8, 9], which has been very problematic throughout the COVID-19 pandemic in particular [10].
@@ 2 The multigroup aNIMFA model
where $y_{i}$ is the local prevalence in group $i$, $z_{ij}$ is the link density between group $j$ and $i$ (which is not necessarily equal to $z_{ji}$), $\delta_{i}$ is the curing rate of group $i$, $\beta_{ij}$ is the infection rate from group $j$ to group $i$, $\zeta_{ij}$ is the link-breaking rate between group $j$ and $i$, $\xi_{ij}$ is the link-creation rate between group $j$ and $i$ and $\bar{y}$ is the global prevalence (or simply the prevalence) of the infectious disease in the whole population, i.e. the average prevalence over all communities. We are implicitly assuming here that all communities have the same size. In general, the functional responses $f_{\textnormal{br,ij}}$ and $f_{\textnormal{cr,ij}}$ can be different between each couple of communities. In particular, when $i=j$, the functional responses $f_{\textnormal{br,ii}}$ and $f_{\textnormal{cr,ii}}$ describe the different responses of each community $i$ to the internal spread of the infectious disease. We represent the exemplifying flow within and between two connected communities in Figure 1.
[…]
Figure 1: Interaction within and between two connected communities $i$ and $j$. Solid lines: change of state of individuals within one community; dashed lines: inter-community infections. Notice that the position of the Susceptible and Infected compartments are switched from the top and the bottom row.
We remark that we only need to keep track of the fraction of Infected individuals in each community, since Susceptible individuals are exactly $1-y_{i}$ at all times, as we are considering an SIS model with no demography spreading in a network of communities with normalized populations.
The $n$ communities can be represented as $n$ nodes in a graph $G$, where each node represents one community. The links in the graph $G$ constitute the connections $\beta_{ji}$ from community $i$ to community $j$. We assume the graph $G$ is strongly connected (i.e. the adjacency matrix $B=(\beta_{ij})$ is irreducible). Although the infection rate $\beta_{ij}$ may be positive from community $j$ to $i$, the adaptive weight $z_{ij}$ may become (temporarily) zero due to the link-adaptivity process. On the other hand, a non-existent link $\beta_{ij}=0$ in the adjacency matrix $B$ can never yield a time-varying (positive) link weight $z_{ij}>0$, because there is simply no interaction from community $j$ to $i$. Consequently, one may chose to ignore entirely the ODEs corresponding to $z_{ij}$ for which $\beta_{ij}=0$, or set the corresponding adaptivity functions and parameters to zero.
@@
From a biological point of view, a natural assumption would be to consider all $\delta_{i}$ equal, since they represent the inverse of the recovery rate of the same disease. However, each region may have a different healthcare quality, leading to potentially slightly different recovery rates $\delta_{i}$. Therefore, we keep the model (2) as general as possible.
@@
Equation (2) is coined the multigroup adaptive N-Intertwined Mean-Field Approximation (multigroup aNIMFA), named after the NIMFA model [27]. On a technical note, Eq. (2) contains at most $n+n^{2}$ equations (one for each community, and at most $n^{2}$ for the connections among them), not $n$, but we will adhere to the definition (2) due its analogy to the static NIMFA model [27]. We assume the parameters $\delta_{i}$, $\beta_{ij}$, $\zeta_{ij}$, $\xi_{ij}$ can be different for each node and link, and are assumed to be non-negative. The standard NIMFA model with link weights $z_{ij}=z_{ij}(0)$ is recovered if $\zeta_{ij}=\xi_{ij}=0$ for all $i$ and $j$. Furthermore, $f_{\textnormal{br,ij}}$ and $f_{\textnormal{cr,ij}}$ are assumed to be non-negative for all $0\leq y\leq 1$ and all $i$ and $j$. We emphasise here that the link density in the aNIMFA model is not necessarily symmetric, i.e. $z_{ij}\neq z_{ji}$. This allows us to capture unilateral decisions to limit or encourage movement between community $i$ and community $j$, but not vice versa. For example, we can think of some country A which lifts restrictions on incoming flights from another country B, even though country B still discourages its inhabitants to travel.
@@
We remark that system (2) may be derived as a mean field limit of a stochastic model, as was done in [28, 29]. In particular, it is a generalization of [29, Eq. (4)], which allows us to include non-local information on the prevalence in the adaptivity mechanism.
@@ 3.3 Basic Reproduction Number
Notice that the matrix $M$ represents infections, whereas the matrix $V$ represents recovery. Then, the basic reproduction number $R_{0}$ follows as the largest eigenvalue $\rho$ of the matrix $F$
@@
If instead we assume no link-breaking at zero prevalence, i.e. $f_{\textnormal{br,ij}}(\mathbf{0})=0$ for all $i,j$, then the steady-state link density $z_{ij}^{\textnormal{DFE}}=1$ for all $i,j$ and the matrix
@@ 3.4 Stability of the DFE
The proof of Corollary 4 coincides, up to extremely minor adjustements, with the second half of the proof of Theorem 1 in [30], which in turn makes use of a few smaller results presented in the same paper; for the sake of brevity, we do not repeat it here.
@@ 3.5 Endemic Equilibria
We now prove, under slightly more restrictive assumptions than $R_{0}>1$, the existence of (at least one) Endemic Equilibrium (EE), in which the prevalence in each community is strictly positive, i.e. $y_{i}>0$ for $i=1,2,\dots,n$. Indeed, assuming that at equilibrium one $y_{i}>0$ necessarily means that all adjacent communities must have strictly positive prevalence in equilibrium. Since the network is assumed to be strongly connected, this implies that all $y_{i}$ are strictly positive at the EE. We remark that, in all generality, we can not prove stability nor instability of the EE(s). We do not prove uniqueness and, as we shall see in Section 4, system (2) admits limit cycles even in a network with only two communities.
@@
We remark, moreover, that the number of EE heavily depends on the specific choices for the function $f_{\textnormal{cr,ij}},f_{\textnormal{br,ij}}$. In order to reach stronger conclusions on their cardinality and stability, one needs to either select specific examples or impose additional conditions.
@@ 4 Numerical simulations
The complexity of the governing Equations (2) complicates deriving further results on the dynamics. We therefore resort to numerical simulations of system (2) in specific scenarios. The simulations in this section have been executed in Matlab and are based on a simple Forward Euler scheme with time step $\Delta t=0.01$ of system (2). In the simulations, we were careful to account for the fact that the adaptivity of the weights $z_{ij}$ may vary the intensity of contacts between two communities connected by the corresponding $\beta_{ij}$, but does not create new connections not encoded in the adjacency matrix $B$. This consideration ensures that the numerical implementation is efficient, as it reduces the number of differential equations that need to be integrated. The parameters $\delta_{i}$, $\beta_{ij}$, $\zeta_{ij}$ and $\xi_{ij}$ are all rates, i.e. $1$ over unit time.
We present three case studies on the multigroup aNIMFA model, each discussing a key aspect of the interplay between infectious disease dynamics and disease awareness.
The codes used to produce the simulations contained in this section are available in the GitHub repository [^1] https://github.com/SaraSottile/AdaptativeMeanField. The parameters for Case Study 1 are provided in Table 1, while the parameters for Case Studies 2 and 3 can be generated from the codes by using the provided random seed. All the Figures are based on a single realization of the code. Refer to Section 6 for the codes used to produce all the Figures.
@@ 4.1 Case study 1: Periodic behaviour
Figure 2: Case study 1: effect of asymmetric response between two communities. Parameters are listed in Table 1. (a) local and average prevalence; (b) link densities. We observe a quick convergence towards a stable limit cycle, representing endemicity of the disease alternating between high and low prevalence in the population.
Parameter Value
$n$ $2$
$\beta_{ij}$ $\left(\begin{array}[]{cc}0.52&0.66\\ 0.03&0.52\end{array}\right)$
$\delta_{1}=\delta_{2}$ $0.50$
$\zeta_{ij}$ $\left(\begin{array}[]{cc}0.42&0.20\\ 0.33&0.62\end{array}\right)$
$\xi_{ij}$ $\left(\begin{array}[]{cc}0.30&0.62\\ 0.27&0.53\end{array}\right)$
$z_{ij}(0)$ $\left(\begin{array}[]{cc}1&1\\ 1&1\end{array}\right)$
$y(0)$ $(0.2,0)$
Table 1: Parameters used for Case study 1, see Figure 2. These values were obtained by sampling, respectively: $\beta_{ij}\sim U([0,1.2]),\delta_{i}=U([0,0.5]),\zeta_{ij}\sim U([0,1]),\xi_{ij}\sim U([0,1])$ for $i,j=1,2$, where $U$ denotes the uniform distribution.
@@
Even though the provided functional responses contain indicator functions, which are not $\mathcal{C}^{1}$-functions over the interval $[0,1]$, we emphasise that other choices of $f_{\text{cr,ii}}$, $f_{\text{br,ii}}$, $f_{\text{cr,ij}}$ and $f_{\text{br,ij}}$ can also lead to periodic behaviour. The benefit of having $\mathcal{C}^{1}$-functions is that the solution is guaranteed to be unique. One method is to approximate the indicator function, for which we verified that periodic solutions also occur. We refer, for a few options of smooth approximations of the indicator function, to Chapter 5.3 in [33].
@@
4.2 Case study 2: Internal and external connections